OpenStudy (anonymous):

Solve: 6x2 – 15x = 0

OpenStudy (bahrom7893):

6x^2 - 15x = 0, pull out an x x(6x - 15) = 0 x = 0 or 6x -15 = 0 x = 0 or 6x = 15 x = 0 or x = 15/6 = 5/2 = 2.5